# A Convex Hull Cheapest Insertion Heuristic for the Non-Euclidean TSP

### v3

@@ 5.2 Performance Analysis
t70 1.27 $(\mathcal{L}_{1}norm)$ 8.6e+02 1.00e+03 9.32e+02 8.92e+02 (1.77e+01) 9.26e+02 (3.58e+01)
1.21 $(4~{}separators)$ 7.7e+02 8.86e+02 8.88e+02 8.17e+02 (6.81e+01) 7.86e+02 (3.68e+01)
1.44 $(16~{}separators)$ 1.2e+03 1.23e+03 1.28e+03 1.19e+03 (1.85e+00) 1.18e+03 (5.96e+00)
1.60 $(64~{}separators)$ 2.1e+03 2.22e+03 2.18e+03 2.06e+03 (4.62e+00) 2.16e+03 (3.96e+01)
rat99 1.24 $(\mathcal{L}_{1}norm)$ 1.6e+03 1.84e+03 1.86e+03 1.65e+03 (6.65e+01) 1.70e+03 (5.32e+01)
1.15 $(4~{}separators)$ 1.4e+03 1.57e+03 1.55e+03 1.40e+03 (1.16e+02) 1.49e+03 (8.25e+01)
1.38 $(16~{}separators)$ 2.1e+03 2.35e+03 2.35e+03 2.09e+03 (7.87e+01) 2.09e+03 (6.16e+01)
1.54 $(64~{}separators)$ 4.0e+03 4.20e+03 4.28e+03 4.03e+03 (3.50e+00) 4.23e+03 (4.25e+01)
roE100 1.25 $(\mathcal{L}_{1}norm)$ 2.9e+04 3.44e+04 3.20e+04 3.01e+04 (1.92e+03) 3.09e+04 (3.76e+02)
1.19 $(4~{}separators)$ 2.4e+04 2.79e+04 2.88e+04 2.59e+04 (4.81e+03) 2.61e+04 (7.11e+02)
1.39 $(16~{}separators)$ 3.7e+04 3.88e+04 4.20e+04 3.71e+04 (8.00e+01) 3.67e+04 (1.49e+02)
1.61 $(64~{}separators)$ 6.5e+04 7.13e+04 7.11e+04 6.59e+04 (1.34e+03) 7.07e+04 (1.28e+03)
pr107 1.17 $(\mathcal{L}_{1}norm)$ 5.2e+04 5.79e+04 5.78e+04 5.28e+04 (1.77e+03) 5.16e+04 (2.58e+02)
1.03 $(4~{}separators)$ 4.6e+04 5.39e+04 5.19e+04 4.91e+04 (4.20e+03) 4.93e+04 (1.26e+03)
1.23 $(16~{}separators)$ 5.4e+04 6.26e+04 6.23e+04 5.58e+04 (2.55e+03) 5.83e+04 (4.51e+02)
1.47 $(64~{}separators)$ 8.2e+04 8.86e+04 8.66e+04 8.43e+04 (4.62e+02) 8.74e+04 (6.62e+02)
bier127 1.28 $(\mathcal{L}_{1}norm)$ 1.6e+05 1.80e+05 1.81e+05 1.68e+05 (7.68e+03) 1.67e+05 (2.63e+03)
1.16 $(4~{}separators)$ 1.4e+05 1.76e+05 1.54e+05 1.42e+05 (7.20e+03) 1.59e+05 (5.69e+03)
1.31 $(16~{}separators)$ 1.9e+05 2.06e+05 2.11e+05 1.92e+05 (1.68e+04) 2.11e+05 (5.72e+03)
1.45 $(64~{}separators)$ 3.8e+05 4.27e+05 3.93e+05 3.85e+05 (6.03e+03) 4.35e+05 (5.68e+03)
pr136 1.25 $(\mathcal{L}_{1}norm)$ 1.3e+05 1.20e+05 1.34e+05 1.31e+05 (4.01e+03) 1.33e+05 (2.35e+03)
1.19 $(4~{}separators)$ 1.2e+05 1.38e+05 1.24e+05 1.21e+05 (9.21e+03) 1.30e+05 (6.21e+03)
1.44 $(16~{}separators)$ 1.6e+05 1.63e+05 1.70e+05 1.61e+05 (2.02e+03) 1.64e+05 (2.56e+03)
1.66 $(64~{}separators)$ 3.0e+05 2.98e+05 3.24e+05 3.15e+05 (2.24e+04) 3.21e+05 (4.35e+03)
d198 1.27 $(\mathcal{L}_{1}norm)$ 2.1e+04 2.35e+04 2.37e+04 1.82e+05 (1.70e+04) 2.62e+04 (3.96e+03)
1.20 $(4~{}separators)$ 1.8e+04 2.10e+04 1.91e+04 1.65e+05 (1.48e+04) 2.20e+04 (2.44e+03)
1.45 $(16~{}separators)$ 2.2e+04 2.41e+04 2.26e+04 1.98e+05 (7.94e+03) 2.66e+04 (4.00e+03)
1.65 $(64~{}separators)$ 4.7e+04 4.35e+04 5.13e+04 2.18e+05 (1.04e+04) 5.24e+04 (1.02e+04)
gil262 1.27 $(\mathcal{L}_{1}norm)$ 3.2e+03 3.67e+03 3.51e+03 1.31e+04 (3.14e+03) 3.65e+03 (1.77e+02)
1.22 $(4~{}separators)$ 2.9e+03 3.20e+03 3.05e+03 1.27e+04 (2.69e+03) 3.17e+03 (5.65e+01)
1.50 $(16~{}separators)$ 3.7e+03 3.97e+03 4.05e+03 1.81e+04 (2.34e+03) 4.01e+03 (1.34e+02)
1.70 $(64~{}separators)$ 8.6e+03 8.53e+03 9.28e+03 2.31e+04 (1.70e+03) 9.02e+03 (9.49e+01)
pr264 1.19 $(\mathcal{L}_{1}norm)$ 1.5e+04 1.83e+04 1.79e+04 3.46e+05 (3.91e+04) 1.85e+04 (1.66e+03)
1.04 $(4~{}separators)$ 1.3e+04 1.63e+04 1.45e+04 3.13e+05 (3.04e+04) 1.62e+04 (1.44e+03)
1.22 $(16~{}separators)$ 2.0e+04 2.17e+04 2.21e+04 3.42e+05 (4.95e+04) 2.25e+04 (3.24e+03)
1.50 $(64~{}separators)$ 3.5e+04 3.81e+04 3.66e+04 3.79e+05 (3.75e+04) 4.06e+04 (3.42e+03)
in318 1.26 $(\mathcal{L}_{1}norm)$ 5.1e+04 5.61e+04 5.40e+04 6.62e+05 (2.59e+04) -
1.19 $(4~{}separators)$ 4.3e+04 4.98e+04 4.67e+04 6.00e+05 (1.76e+04) -
1.46 $(16~{}separators)$ 5.1e+04 5.91e+04 5.40e+04 7.15e+05 (4.03e+04) -
1.68 $(64~{}separators)$ 1.2e+05 1.09e+05 1.22e+05 7.75e+05 (1.98e+04) -
rd400 1.27 $(\mathcal{L}_{1}norm)$ 3.5e+05 4.02e+05 3.78e+05 7.33e+06 (1.51e+05) -
1.23 $(4~{}separators)$ 2.9e+05 3.48e+05 3.13e+05 6.81e+06 (1.80e+05) -
1.53 $(16~{}separators)$ 3.5e+05 4.03e+05 3.59e+05 7.85e+06 (1.81e+05) -
[…]

### v5

@@ 5.2 Performance Analysis
eil51 1.28 $(\mathcal{L}_{1}norm)$ 5.76e+02 6.62e+02 6.16e+02 5.75e+02 (1.82e+00) 5.69e+02 (8.36e+00)
1.18 $(4~{}separators)$ 4.83e+02 5.60e+02 5.54e+02 4.90e+02 (1.44e+01) 4.86e+02 (1.05e+01)
1.42 $(16~{}separators)$ 7.33e+02 7.91e+02 8.20e+02 7.30e+02 (9.95e+00) 7.40e+02 (1.00e+01)
1.55 $(64~{}separators)$ 1.31e+03 1.35e+03 1.33e+03 1.31e+03 (2.20e+00) 1.37e+03 (1.34e+01)
erlin52 1.27 $(\mathcal{L}_{1}norm)$ 9.92e+03 1.12e+04 1.08e+04 9.93e+03 (2.32e+02) 9.89e+03 (4.01e+01)
1.14 $(4~{}separators)$ 8.98e+03 9.96e+03 9.81e+03 8.97e+03 (6.06e+00) 9.02e+03 (5.98e+01)
1.30 $(16~{}separators)$ 1.37e+04 1.55e+04 1.50e+04 1.36e+04 (3.30e+01) 1.41e+04 (1.94e+02)
1.43 $(64~{}separators)$ 2.16e+04 2.33e+04 2.25e+04 2.16e+04 (2.73e+01) 2.28e+04 (5.32e+01)
t70 1.27 $(\mathcal{L}_{1}norm)$ 8.60e+02 1.00e+03 9.32e+02 8.92e+02 (1.77e+01) 9.26e+02 (3.58e+01)
1.21 $(4~{}separators)$ 7.69e+02 8.86e+02 8.88e+02 8.17e+02 (6.81e+01) 7.86e+02 (3.68e+01)
1.44 $(16~{}separators)$ 1.19e+03 1.23e+03 1.28e+03 1.19e+03 (1.85e+00) 1.18e+03 (5.96e+00)
1.60 $(64~{}separators)$ 2.06e+03 2.22e+03 2.18e+03 2.06e+03 (4.62e+00) 2.16e+03 (3.96e+01)
eil76 1.28 $(\mathcal{L}_{1}norm)$ 7.50e+02 8.74e+02 7.82e+02 7.50e+02 (3.51e+00) 7.57e+02 (8.30e+00)
1.19 $(4~{}separators)$ 6.27e+02 7.05e+02 6.99e+02 6.23e+02 (5.26e+00) 6.34e+02 (3.23e+01)
1.40 $(16~{}separators)$ 9.17e+02 9.48e+02 9.75e+02 9.12e+02 (4.06e+00) 8.88e+02 (1.24e+01)
1.52 $(64~{}separators)$ 1.61e+03 1.69e+03 1.69e+03 1.61e+03 (1.47e+00) 1.72e+03 (4.17e+01)
pr76 1.26 $(\mathcal{L}_{1}norm)$ 1.46e+05 1.78e+05 1.55e+05 1.45e+05 (6.62e+02) 1.44e+05 (8.55e+02)
1.19 $(4~{}separators)$ 1.22e+05 1.53e+05 1.38e+05 1.30e+05 (1.32e+04) 1.22e+05 (1.33e+03)
1.46 $(16~{}separators)$ 1.85e+05 1.99e+05 2.19e+05 1.85e+05 (2.20e+02) 1.92e+05 (2.33e+03)
1.62 $(64~{}separators)$ 3.72e+05 3.97e+05 3.76e+05 3.72e+05 (3.47e+02) 4.00e+05 (3.13e+03)
rat99 1.24 $(\mathcal{L}_{1}norm)$ 1.63e+03 1.84e+03 1.86e+03 1.65e+03 (6.65e+01) 1.70e+03 (5.32e+01)
1.15 $(4~{}separators)$ 1.36e+03 1.57e+03 1.55e+03 1.40e+03 (1.16e+02) 1.49e+03 (8.25e+01)
1.38 $(16~{}separators)$ 2.08e+03 2.35e+03 2.35e+03 2.09e+03 (7.87e+01) 2.09e+03 (6.16e+01)
1.54 $(64~{}separators)$ 4.03e+03 4.20e+03 4.28e+03 4.03e+03 (3.50e+00) 4.23e+03 (4.25e+01)
roC100 1.25 $(\mathcal{L}_{1}norm)$ 2.73e+04 3.13e+04 3.17e+04 2.97e+04 (2.39e+03) 3.04e+04 (6.06e+02)
1.17 $(4~{}separators)$ 2.36e+04 2.87e+04 2.82e+04 2.56e+04 (3.33e+03) 2.63e+04 (1.68e+03)
1.41 $(16~{}separators)$ 3.74e+04 4.04e+04 4.19e+04 3.74e+04 (6.68e+01) 3.81e+04 (5.33e+02)
1.61 $(64~{}separators)$ 7.14e+04 7.83e+04 7.48e+04 7.21e+04 (1.63e+03) 7.72e+04 (5.90e+01)
roD100 1.25 $(\mathcal{L}_{1}norm)$ 2.69e+04 3.26e+04 3.16e+04 2.85e+04 (7.79e+02) 3.02e+04 (8.12e+02)
1.18 $(4~{}separators)$ 2.36e+04 2.60e+04 2.77e+04 2.47e+04 (1.03e+03) 2.58e+04 (2.34e+02)
1.37 $(16~{}separators)$ 3.39e+04 3.87e+04 4.09e+04 3.37e+04 (3.37e+02) 3.59e+04 (1.10e+03)
1.56 $(64~{}separators)$ 6.66e+04 7.17e+04 7.33e+04 6.66e+04 (1.57e+03) 7.08e+04 (1.36e+03)
roE100 1.25 $(\mathcal{L}_{1}norm)$ 2.87e+04 3.44e+04 3.20e+04 3.01e+04 (1.92e+03) 3.09e+04 (3.76e+02)
1.19 $(4~{}separators)$ 2.36e+04 2.79e+04 2.88e+04 2.59e+04 (4.81e+03) 2.61e+04 (7.11e+02)
1.39 $(16~{}separators)$ 3.72e+04 3.88e+04 4.20e+04 3.71e+04 (8.00e+01) 3.67e+04 (1.49e+02)
1.61 $(64~{}separators)$ 6.53e+04 7.13e+04 7.11e+04 6.59e+04 (1.34e+03) 7.07e+04 (1.28e+03)
roA100 1.24 $(\mathcal{L}_{1}norm)$ 2.73e+04 3.17e+04 3.26e+04 2.86e+04 (8.06e+02) 3.00e+04 (1.50e+03)
1.15 $(4~{}separators)$ 2.37e+04 2.84e+04 2.79e+04 2.41e+04 (1.42e+03) 2.56e+04 (6.70e+01)
1.37 $(16~{}separators)$ 3.65e+04 3.94e+04 4.13e+04 3.65e+04 (3.14e+02) 3.70e+04 (5.09e+02)
1.56 $(64~{}separators)$ 7.22e+04 7.45e+04 7.52e+04 7.21e+04 (8.18e+01) 7.64e+04 (4.39e+02)
roB100 1.25 $(\mathcal{L}_{1}norm)$ 2.95e+04 3.57e+04 3.23e+04 3.08e+04 (2.27e+03) 3.14e+04 (8.75e+02)
1.18 $(4~{}separators)$ 2.68e+04 2.98e+04 2.85e+04 2.68e+04 (2.65e+02) 2.90e+04 (3.14e+02)
1.41 $(16~{}separators)$ 3.81e+04 4.02e+04 4.31e+04 3.84e+04 (1.22e+03) 3.80e+04 (2.09e+02)
1.62 $(64~{}separators)$ 7.68e+04 8.02e+04 8.20e+04 7.66e+04 (1.64e+02) 8.10e+04 (1.72e+03)
rd100 1.27 $(\mathcal{L}_{1}norm)$ 1.07e+04 1.23e+04 1.18e+04 1.07e+04 (1.10e+01) 1.11e+04 (4.72e+02)
1.21 $(4~{}separators)$ 9.02e+03 9.98e+03 1.06e+04 9.34e+03 (3.43e+02) 9.79e+03 (4.19e+01)
1.48 $(16~{}separators)$ 1.36e+04 1.34e+04 1.61e+04 1.35e+04 (2.09e+01) 1.37e+04 (2.93e+02)
1.70 $(64~{}separators)$ 2.86e+04 2.99e+04 2.95e+04 2.86e+04 (2.59e+01) 3.06e+04 (5.62e+02)
eil101 1.28 $(\mathcal{L}_{1}norm)$ 8.82e+02 9.76e+02 9.10e+02 8.82e+02 (4.21e+00) 9.50e+02 (1.53e+01)
1.21 $(4~{}separators)$ 7.22e+02 8.65e+02 7.59e+02 7.34e+02 (3.06e+01) 7.83e+02 (3.06e+01)
1.48 $(16~{}separators)$ 1.02e+03 1.11e+03 1.16e+03 1.03e+03 (2.30e+01) 1.03e+03 (4.35e+01)
1.68 $(64~{}separators)$ 2.07e+03 2.13e+03 2.18e+03 2.06e+03 (1.67e+00) 2.20e+03 (1.78e+01)
in105 1.23 $(\mathcal{L}_{1}norm)$ 1.71e+04 1.94e+04 2.10e+04 1.84e+04 (1.61e+03) 1.80e+04 (5.70e+02)
1.08 $(4~{}separators)$ 1.68e+04 1.95e+04 1.86e+04 1.69e+04 (9.56e+02) 1.82e+04 (3.75e+02)
1.38 $(16~{}separators)$ 2.34e+04 2.55e+04 2.79e+04 2.34e+04 (5.82e+01) 2.46e+04 (2.08e+00)
1.63 $(64~{}separators)$ 4.25e+04 4.41e+04 4.73e+04 4.22e+04 (2.26e+02) 4.45e+04 (7.86e+02)
Table 2 (contd.): TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
pr107 1.17 $(\mathcal{L}_{1}norm)$ 5.18e+04 5.79e+04 5.78e+04 5.28e+04 (1.77e+03) 5.16e+04 (2.58e+02)
1.03 $(4~{}separators)$ 4.64e+04 5.39e+04 5.19e+04 4.91e+04 (4.20e+03) 4.93e+04 (1.26e+03)
1.23 $(16~{}separators)$ 5.41e+04 6.26e+04 6.23e+04 5.58e+04 (2.55e+03) 5.83e+04 (4.51e+02)
1.47 $(64~{}separators)$ 8.22e+04 8.86e+04 8.66e+04 8.43e+04 (4.62e+02) 8.74e+04 (6.62e+02)
pr124 1.26 $(\mathcal{L}_{1}norm)$ 7.14e+04 8.60e+04 8.36e+04 8.05e+04 (6.22e+03) 7.26e+04 (1.91e+03)
1.20 $(4~{}separators)$ 6.85e+04 8.66e+04 8.38e+04 7.13e+04 (6.40e+03) 7.37e+04 (2.45e+03)
1.48 $(16~{}separators)$ 1.10e+05 1.12e+05 1.20e+05 1.10e+05 (1.87e+02) 1.10e+05 (2.33e+03)
1.77 $(64~{}separators)$ 2.30e+05 2.41e+05 2.41e+05 2.34e+05 (6.89e+03) 2.52e+05 (1.48e+03)
bier127 1.28 $(\mathcal{L}_{1}norm)$ 1.60e+05 1.80e+05 1.81e+05 1.68e+05 (7.68e+03) 1.67e+05 (2.63e+03)
1.16 $(4~{}separators)$ 1.42e+05 1.76e+05 1.54e+05 1.42e+05 (7.20e+03) 1.59e+05 (5.69e+03)
1.31 $(16~{}separators)$ 1.86e+05 2.06e+05 2.11e+05 1.92e+05 (1.68e+04) 2.11e+05 (5.72e+03)
1.45 $(64~{}separators)$ 3.80e+05 4.27e+05 3.93e+05 3.85e+05 (6.03e+03) 4.35e+05 (5.68e+03)
h130 1.28 $(\mathcal{L}_{1}norm)$ 8.14e+03 1.01e+04 9.28e+03 8.50e+03 (4.27e+02) 8.79e+03 (1.36e+02)
1.19 $(4~{}separators)$ 6.90e+03 8.01e+03 7.64e+03 7.13e+03 (7.22e+02) 7.37e+03 (1.58e+02)
1.45 $(16~{}separators)$ 9.99e+03 9.98e+03 1.13e+04 1.00e+04 (3.65e+02) 9.99e+03 (8.62e+01)
1.65 $(64~{}separators)$ 2.13e+04 2.18e+04 2.26e+04 2.21e+04 (1.20e+03) 2.26e+04 (2.55e+02)
pr136 1.25 $(\mathcal{L}_{1}norm)$ 1.29e+05 1.20e+05 1.34e+05 1.31e+05 (4.01e+03) 1.33e+05 (2.35e+03)
1.19 $(4~{}separators)$ 1.17e+05 1.38e+05 1.24e+05 1.21e+05 (9.21e+03) 1.30e+05 (6.21e+03)
1.44 $(16~{}separators)$ 1.61e+05 1.63e+05 1.70e+05 1.61e+05 (2.02e+03) 1.64e+05 (2.56e+03)
1.66 $(64~{}separators)$ 3.02e+05 2.98e+05 3.24e+05 3.15e+05 (2.24e+04) 3.21e+05 (4.35e+03)
pr144 1.24 $(\mathcal{L}_{1}norm)$ 7.36e+04 7.66e+04 8.13e+04 7.87e+04 (8.92e+03) 7.27e+04 (2.88e+02)
1.17 $(4~{}separators)$ 6.82e+04 6.83e+04 7.98e+04 7.39e+04 (9.32e+03) 6.99e+04 (7.85e+02)
1.33 $(16~{}separators)$ 9.43e+04 1.02e+05 1.10e+05 9.88e+04 (8.50e+03) 1.00e+05 (4.33e+03)
1.57 $(64~{}separators)$ 1.78e+05 1.87e+05 1.90e+05 1.87e+05 (8.40e+03) 1.91e+05 (2.86e+03)
roB150 1.25 $(\mathcal{L}_{1}norm)$ 3.53e+04 4.19e+04 3.83e+04 3.97e+04 (6.10e+03) 3.96e+04 (3.49e+02)
1.17 $(4~{}separators)$ 3.03e+04 3.52e+04 3.38e+04 3.39e+04 (4.07e+03) 3.38e+04 (1.01e+03)
1.40 $(16~{}separators)$ 4.35e+04 4.61e+04 4.76e+04 4.66e+04 (3.44e+03) 4.49e+04 (8.14e+02)
1.62 $(64~{}separators)$ 8.31e+04 8.81e+04 8.94e+04 9.73e+04 (1.66e+04) 8.77e+04 (1.55e+03)
roA150 1.25 $(\mathcal{L}_{1}norm)$ 3.58e+04 4.23e+04 3.94e+04 4.03e+04 (2.29e+03) 3.95e+04 (1.17e+03)
1.17 $(4~{}separators)$ 2.93e+04 3.69e+04 3.27e+04 3.28e+04 (2.69e+03) 3.35e+04 (1.93e+03)
1.39 $(16~{}separators)$ 4.16e+04 4.44e+04 4.55e+04 4.50e+04 (7.58e+03) 4.39e+04 (1.07e+03)
1.59 $(64~{}separators)$ 8.92e+04 9.50e+04 9.36e+04 9.70e+04 (1.03e+04) 9.53e+04 (1.62e+03)
h150 1.27 $(\mathcal{L}_{1}norm)$ 8.71e+03 9.93e+03 1.01e+04 1.02e+04 (9.87e+02) 9.74e+03 (3.49e+02)
1.16 $(4~{}separators)$ 7.05e+03 8.79e+03 8.37e+03 8.31e+03 (8.87e+02) 7.80e+03 (7.46e+02)
1.35 $(16~{}separators)$ 1.06e+04 1.00e+04 1.14e+04 1.19e+04 (1.98e+03) 1.05e+04 (2.90e+01)
1.50 $(64~{}separators)$ 2.00e+04 2.13e+04 2.13e+04 2.24e+04 (1.94e+03) 2.17e+04 (4.61e+02)
pr152 1.22 $(\mathcal{L}_{1}norm)$ 9.20e+04 9.64e+04 1.05e+05 9.58e+04 (1.34e+04) 9.70e+04 (3.23e+03)
1.09 $(4~{}separators)$ 7.93e+04 9.85e+04 9.22e+04 8.54e+04 (1.96e+03) 8.66e+04 (4.66e+03)
1.31 $(16~{}separators)$ 1.10e+05 1.32e+05 1.21e+05 1.13e+05 (6.32e+03) 1.18e+05 (2.52e+03)
1.46 $(64~{}separators)$ 1.82e+05 1.90e+05 1.92e+05 1.92e+05 (1.06e+04) 1.92e+05 (9.75e+02)
u159 1.24 $(\mathcal{L}_{1}norm)$ 5.58e+04 5.96e+04 6.16e+04 6.45e+04 (1.25e+04) 6.00e+04 (1.81e+03)
1.13 $(4~{}separators)$ 5.09e+04 5.49e+04 5.67e+04 5.75e+04 (7.10e+03) 5.60e+04 (1.54e+03)
1.40 $(16~{}separators)$ 7.22e+04 7.33e+04 7.79e+04 8.02e+04 (8.72e+03) 7.35e+04 (4.83e+03)
1.63 $(64~{}separators)$ 1.38e+05 1.37e+05 1.43e+05 1.56e+05 (1.95e+04) 1.42e+05 (7.34e+02)
rat195 1.24 $(\mathcal{L}_{1}norm)$ 3.31e+03 3.62e+03 3.52e+03 5.54e+03 (9.97e+02) 3.47e+03 (1.00e+02)
1.15 $(4~{}separators)$ 2.67e+03 3.10e+03 2.92e+03 4.57e+03 (6.53e+02) 3.04e+03 (7.75e+00)
1.38 $(16~{}separators)$ 3.45e+03 3.68e+03 3.74e+03 7.29e+03 (1.09e+03) 3.73e+03 (8.31e+01)
1.54 $(64~{}separators)$ 7.77e+03 7.67e+03 8.31e+03 1.16e+04 (9.04e+02) 7.93e+03 (8.93e+01)
d198 1.27 $(\mathcal{L}_{1}norm)$ 2.00e+04 2.24e+04 2.14e+04 2.86e+04 (7.41e+03) 2.18e+04 (7.32e+02)
1.11 $(4~{}separators)$ 1.77e+04 1.84e+04 1.85e+04 2.41e+04 (4.36e+03) 1.90e+04 (9.05e+01)
1.51 $(16~{}separators)$ 2.05e+04 2.33e+04 2.21e+04 3.02e+04 (5.36e+03) 2.29e+04 (6.54e+02)
1.99 $(64~{}separators)$ 4.13e+04 4.28e+04 4.26e+04 5.82e+04 (1.53e+04) 4.30e+04 (4.43e+02)
roA200 1.25 $(\mathcal{L}_{1}norm)$ 4.06e+04 4.53e+04 4.48e+04 7.97e+04 (2.14e+04) 4.44e+04 (1.47e+03)
1.18 $(4~{}separators)$ 3.45e+04 4.18e+04 3.71e+04 6.97e+04 (1.40e+04) 3.93e+04 (6.54e+02)
1.42 $(16~{}separators)$ 4.60e+04 5.28e+04 5.12e+04 1.05e+05 (1.58e+04) 4.96e+04 (7.63e+02)
1.62 $(64~{}separators)$ 9.66e+04 9.82e+04 1.07e+05 1.68e+05 (1.59e+04) 1.02e+05 (9.58e+02)
Table 2 (contd.): TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
roB200 1.25 $(\mathcal{L}_{1}norm)$ 4.01e+04 4.77e+04 4.55e+04 8.87e+04 (1.35e+04) 4.52e+04 (3.75e+02)
1.17 $(4~{}separators)$ 3.41e+04 3.80e+04 3.80e+04 6.56e+04 (1.19e+04) 3.80e+04 (2.01e+03)
1.40 $(16~{}separators)$ 4.67e+04 5.47e+04 5.12e+04 9.97e+04 (1.99e+04) 5.13e+04 (1.31e+03)
1.60 $(64~{}separators)$ 1.07e+05 1.08e+05 1.13e+05 1.71e+05 (1.55e+04) 1.11e+05 (1.06e+03)
ts225 1.26 $(\mathcal{L}_{1}norm)$ 1.61e+05 1.60e+05 1.81e+05 5.37e+05 (7.56e+04) 1.57e+05 (3.84e+03)
1.17 $(4~{}separators)$ 1.61e+05 1.57e+05 1.94e+05 5.16e+05 (1.54e+05) 1.61e+05 (1.67e+03)
1.42 $(16~{}separators)$ 3.80e+05 3.84e+05 3.86e+05 8.60e+05 (7.77e+04) 3.96e+05 (1.07e+04)
1.54 $(64~{}separators)$ 6.84e+05 6.69e+05 6.95e+05 1.10e+06 (7.44e+04) 7.72e+05 (1.98e+04)
tsp225 1.26 $(\mathcal{L}_{1}norm)$ 4.55e+03 5.66e+03 4.99e+03 1.31e+04 (2.86e+03) 4.96e+03 (8.74e+01)
1.19 $(4~{}separators)$ 4.48e+03 5.57e+03 4.75e+03 1.30e+04 (3.98e+03) 5.13e+03 (2.15e+02)
1.45 $(16~{}separators)$ 5.80e+03 6.74e+03 6.29e+03 1.81e+04 (2.62e+03) 6.49e+03 (2.41e+02)
1.64 $(64~{}separators)$ 1.40e+04 1.43e+04 1.50e+04 2.85e+04 (4.93e+03) 1.48e+04 (3.46e+02)
pr226 1.23 $(\mathcal{L}_{1}norm)$ 9.61e+04 1.13e+05 1.16e+05 3.69e+05 (1.28e+05) 1.06e+05 (9.36e+03)
1.19 $(4~{}separators)$ 9.13e+04 1.20e+05 1.10e+05 3.84e+05 (9.76e+04) 1.03e+05 (4.39e+03)
1.48 $(16~{}separators)$ 1.44e+05 1.53e+05 1.66e+05 5.73e+05 (1.59e+05) 1.61e+05 (3.55e+03)
1.87 $(64~{}separators)$ 2.85e+05 2.98e+05 3.03e+05 8.84e+05 (2.32e+05) 3.12e+05 (2.45e+03)
gil262 1.27 $(\mathcal{L}_{1}norm)$ 3.17e+03 3.67e+03 3.51e+03 1.31e+04 (3.14e+03) 3.65e+03 (1.77e+02)
1.22 $(4~{}separators)$ 2.87e+03 3.20e+03 3.05e+03 1.27e+04 (2.69e+03) 3.17e+03 (5.65e+01)
1.50 $(16~{}separators)$ 3.66e+03 3.97e+03 4.05e+03 1.81e+04 (2.34e+03) 4.01e+03 (1.34e+02)
1.70 $(64~{}separators)$ 8.58e+03 8.53e+03 9.28e+03 2.31e+04 (1.70e+03) 9.02e+03 (9.49e+01)
pr264 1.19 $(\mathcal{L}_{1}norm)$ 6.12e+04 6.06e+04 6.64e+04 2.64e+05 (7.30e+04) 6.21e+04 (3.30e+03)
1.03 $(4~{}separators)$ 5.48e+04 6.03e+04 6.00e+04 2.16e+05 (5.41e+04) 6.14e+04 (1.80e+03)
1.50 $(16~{}separators)$ 6.51e+04 7.01e+04 7.54e+04 4.05e+05 (1.04e+05) 7.40e+04 (2.16e+03)
2.00 $(64~{}separators)$ 1.18e+05 1.28e+05 1.25e+05 6.65e+05 (1.15e+05) 1.27e+05 (3.80e+03)
pr299 1.23 $(\mathcal{L}_{1}norm)$ 6.40e+04 6.92e+04 6.93e+04 3.86e+05 (4.25e+04) 7.10e+04 (2.54e+03)
1.22 $(4~{}separators)$ 5.48e+04 6.00e+04 6.01e+04 3.85e+05 (8.54e+04) 6.45e+04 (1.61e+03)
1.49 $(16~{}separators)$ 6.82e+04 7.80e+04 7.69e+04 4.56e+05 (9.63e+04) 8.23e+04 (1.16e+03)
1.79 $(64~{}separators)$ 1.64e+05 1.63e+05 1.75e+05 6.20e+05 (4.02e+04) 1.75e+05 (5.08e+03)
in318 1.26 $(\mathcal{L}_{1}norm)$ 5.40e+04 6.00e+04 5.99e+04 3.89e+05 (7.69e+04) 6.07e+04 (1.57e+03)
1.21 $(4~{}separators)$ 4.92e+04 5.57e+04 5.30e+04 3.47e+05 (3.91e+04) 5.64e+04 (7.37e+02)
1.46 $(16~{}separators)$ 6.42e+04 6.96e+04 7.02e+04 4.37e+05 (5.57e+04) 7.11e+04 (2.23e+03)
1.69 $(64~{}separators)$ 1.40e+05 1.44e+05 1.53e+05 5.43e+05 (3.81e+04) 1.52e+05 (2.09e+03)
in318 1.26 $(\mathcal{L}_{1}norm)$ 5.40e+04 6.00e+04 5.99e+04 3.87e+05 (7.16e+04) 6.03e+04 (1.13e+03)
1.21 $(4~{}separators)$ 4.92e+04 5.57e+04 5.30e+04 3.48e+05 (4.80e+04) 5.66e+04 (7.92e+02)
1.46 $(16~{}separators)$ 6.42e+04 6.96e+04 7.02e+04 4.44e+05 (3.70e+04) 7.21e+04 (1.18e+03)
1.69 $(64~{}separators)$ 1.40e+05 1.44e+05 1.53e+05 5.28e+05 (9.64e+04) 1.51e+05 (2.32e+03)
rd400 1.27 $(\mathcal{L}_{1}norm)$ 2.09e+04 2.35e+04 2.37e+04 1.82e+05 (1.70e+04) 2.62e+04 (3.96e+03)
1.20 $(4~{}separators)$ 1.77e+04 2.10e+04 1.91e+04 1.65e+05 (1.48e+04) 2.20e+04 (2.44e+03)
1.45 $(16~{}separators)$ 2.18e+04 2.41e+04 2.26e+04 1.98e+05 (7.94e+03) 2.66e+04 (4.00e+03)
1.65 $(64~{}separators)$ 4.75e+04 4.35e+04 5.13e+04 2.18e+05 (1.04e+04) 5.24e+04 (1.02e+04)
fl417 1.19 $(\mathcal{L}_{1}norm)$ 1.47e+04 1.83e+04 1.79e+04 3.46e+05 (3.91e+04) 1.85e+04 (1.66e+03)
1.04 $(4~{}separators)$ 1.30e+04 1.63e+04 1.45e+04 3.13e+05 (3.04e+04) 1.62e+04 (1.44e+03)
1.22 $(16~{}separators)$ 2.00e+04 2.17e+04 2.21e+04 3.42e+05 (4.95e+04) 2.25e+04 (3.24e+03)
1.50 $(64~{}separators)$ 3.47e+04 3.81e+04 3.66e+04 3.79e+05 (3.75e+04) 4.06e+04 (3.42e+03)
pr439 1.28 $(\mathcal{L}_{1}norm)$ 1.44e+05 1.57e+05 1.56e+05 1.57e+06 (1.40e+05) 1.70e+05 (7.26e+03)
1.29 $(4~{}separators)$ 1.32e+05 1.50e+05 1.44e+05 1.54e+06 (1.43e+05) 1.68e+05 (3.63e+03)
1.60 $(16~{}separators)$ 1.70e+05 1.84e+05 1.83e+05 1.79e+06 (7.26e+04) 2.10e+05 (5.33e+03)
1.99 $(64~{}separators)$ 4.03e+05 4.02e+05 4.19e+05 2.09e+06 (9.10e+04) 4.56e+05 (3.84e+03)
pcb442 1.27 $(\mathcal{L}_{1}norm)$ 6.28e+04 7.58e+04 7.08e+04 6.87e+05 (4.61e+04) 8.13e+04 (1.34e+03)
1.22 $(4~{}separators)$ 5.71e+04 6.57e+04 6.34e+04 6.49e+05 (4.51e+04) 7.72e+04 (2.30e+03)
[…]